OpenStudy (anonymous):

(-5+3i)-(-8+2i) simplify

OpenStudy (triciaal):

remember negative times negative is positive and i^2 =-1

OpenStudy (triciaal):

what do you have?

OpenStudy (agent0smith):

(-5+3i)-(-8+2i) = -5 + 3i + 8 - 2i now just combine like terms.